# library.bib

### v1

@@ A.1 Simulation Setup and Transport Regime
In Section 3, we describe how the two datasets are generated using the subsurface simulation software Pflotran [pflotran], which solves the coupled mass and energy conservation equations. Here, we provide additional technical details and modeling assumptions relevant for reproducibility. Furthermore, we show that for the selected parameters, the heat transport in our system is dominated by advection by a theoretical analysis.
@@
Our data was generated on 12.8km×$\times$12.8km×$\times$5m with a cell size of 5m×$\times$5m×$\times$5m for the baseline simulations. To test scalability, we also simulate a domain that is twice as large in both $x$- and $y$-dimension for the baseline dataset, but only twice as large in $y$-dimension for the more realistic dataset due to dimension restrictions of the available input data.
The baseline dataset uses a constant hydraulic pressure gradient $\nabla p$ of 0.003 [pelzer2024, geologica]; permeability field $k$ is generated using Perlin noise [perlin] within (1.02E-11, 5.10E-9) m2, and $100$ randomly placed heat pumps, which are all operating with a 5 i.e., C injection temperature difference compared to the surrounding groundwater and an injection rate of 0.00024 m3/s. All values are based on realistic parameter ranges in the region of Munich[geologica, geokw22].
For the realistic dataset, instead of random permeability fields $k$, we use permeability field data that is derived from borehole measurements in the Munich region[lfu_gepo]. Furthermore, based on subsurface measurements, we set $\nabla p=0.0025$ for these simulations. All other simulation parameters remain identical to the baseline. Other subsurface and operational parameters are taken directly from [pelzer2024], which also includes additional information about solver setup and boundaries. For mathematical details of the governing equations, we refer the reader to [pflotran_theory, anderson2007introducing, delleur2016elementary].
Simplifications
@@
To quantify whether heat transport in our system is dominated by advection or diffusion, we compute the dimensionless Péclet number [rapp2017peclet], which is defined as
@@
$\alpha=\frac{k}{\rho c_{p}},$
with $k$ the thermal conductivity, $\rho$ the density, and $c_{p}$ the specific heat capacity. We take the parameters of the solid phase of our simulation inputs of
  • Thermal conductivity: $k=0.65\,\mathrm{W/(m\cdot K)}$
@@
[…]
@@
A.2 Neural Networks, Hyperparameters and Training Details
This section provides implementation and training details for all neural network models used in this work. We begin with a description of the baseline UNet architecture, which also forms the foundation of both LGCNN and DDU-Net. Then, we outline the hyperparameter optimization process and present the final configurations per model and dataset. All models are trained using PyTorch[paszke2017automatic].
UNet Architecture
The UNet architecture used in this work is based on the original design presented by Ronneberger et al. [ronneberger2015u], with several modifications to tailor it to the specific needs of our task of predicting temperature fields. In this section, we introduce the key concepts that define the architecture and explain how they contribute to the model’s design. These concepts will be employed in the hyperparameter search to identify the most suitable configuration, taking into account memory and data limitations.
@@ A.2 Neural Networks, Hyperparameters and Training Details
The training process also involves several hyperparameters. The learning rate controls how quickly the model adjusts its parameters during training. The weight decay parameter helps prevent overfitting by penalizing large weights in the model. Furthermore, the Adam optimizer [kingma2014adam] is employed as optimizer. For the realistic permeability field, we additionally introduce the optimizable hyperparameter optimizer switch, which, when enabled, switches the optimizer to LBFGS after 90 epochs.
During inference, each datapoint is processed as a whole, but during training, they are loaded in smaller patches cut out from the datapoint. Optimized hyperparameters include their overlap, i.e., inverse of skip per direction, and their size, called box length. The data split is untouched by this, i.e. one simulation run per train / val / test separately. This is important to not overlap test patches with training ones.
A.2.1 Hyperparameters and Hyperparameter Optimization
We optimize the hyperparameters of our architecture, training process and data loading in several rounds with Optuna [optuna_2019] and additional manual tweaking. Optuna performs optimization using a tree-structured Parzen estimator algorithm. An overlook of all considered hyperparameters, their ranges and our final choice, as well as the hyperparameters fixed during optimization is provided per used model architecture and dataset, e.g., purely data-driven approaches, LGCNN on individual steps or the full pipeline, on datasets of random $k$ versus realistic $k$.
@@ A.2.1 Hyperparameters and Hyperparameter Optimization
The inputs to the model are $pki$ (pressure field, permeability field, and location of heat pumps). For training, the Mean Absolute Error (MAE) loss function was used, and optimized with the Adam optimizer. The learning rate is set to $1\times 10^{-5}$. The model is trained for 10,000 epochs.
Hyperparameters for purely data-driven approaches
The values considered during hyperparameter optimization for the UNet and DDUNet, trained on 73 datapoints and performed using Optuna [optuna_2019], along with the best settings found, are listed in Table 6. Certain hyperparameters were fixed: we set the number of epochs to 750, with an early stopping criterion based on validation loss and a patience of 80 epochs. Additionally, we note that some hyperparameter combinations (e.g., 32 initial features, depth 6, and 3 convolutions per layer with a kernel size of 7) caused memory issues, leading to their exclusion from the hyperparameter search.
@@
Hyperparameter Range $pki\rightarrow v_{x}v_{y}$    (Step 1) $kiv_{x}v_{y}ss_{o}\rightarrow T$    (Step 3) $pki\rightarrow T$    (Full)
@@
After the hyperparameter search, the values corresponding to the best-performing model (based on Huber loss for the validation dataset) were selected. With these values fixed, five models were trained using different randomly sampled initializations to evaluate sensitivity to random initialization, for these values see Tables 11 and 10.
LGCNN Hyperparameters - on Random Permeability
The values considered during hyperparameter optimization with Optuna [optuna_2019] and the best settings found for both steps of LGCNN are listed in Table 7. Although the optimization was originally run for 100 epochs, the optimum was consistently found within the first 25 epochs. Therefore, to reduce computation cost, we therefore conservatively lowered the maximum number of epochs to 50. This adjustment does not affect any of the reported metrics in the paper.
@@
Parameter Range 1st step 3rd step
@@
depth $v:$ 1, 2, 3, 4 4 4
$T:$ 1, 2, 3
Fixed parameters for this hyperparameter search are the learning rate (fixed at $10^{-4}$), ReLU as activation function, the batch size of 20, and the use of a batch normalization layer within the inner blocks of the UNet architeture. The validation loss used for selecting the optimal model is the MAE.
LGCNN Hyperparameters - on Real Permeability
The values considered during hyperparameter optimization on the dataset with a more realistic permeability field were selected using Optuna [optuna_2019], and are summarized in  Table 8, along with the best configurations found for both steps of the LGCNN. The optimization was run for up to 100 epochs. For more background on the network architecture and the various hyperparameters, cf. A.2. Several hyperparameters were fixed during this process. These include a constant learning rate schedule, an Adam optimizer with a weight decay of $10^{-4}$, and, when enabled, a switch to LBFGS after 90 epochs. Fixed architectural parameters include a convolutional stride and dilation of 1. During training, the inputs were cut out from the full datapoints. For model comparison, the validation loss was consistently computed using the Huber loss.
@@
No. initial features 8, 16, 32 32 16
kernel size 3, 5 3 5
depth 4, 5, 6 5 6
@@ A.3 Additional Experimental Results
This section provides additional experimental results. While the main results section focused only on the test and scaling datasets, we also include here the metric values on the training and validation datasets. For completeness and easier comparison, the test and scaling metrics are re-listed as well.
Purely data-driven UNet, DDU-Net: Metrics of training and ablation study
Table 9 presents all the metrics for predicting the temperature field directly from the inputs $pki$ using a data-driven approach, evaluated on the training, validation, and test datasets. The results are provided for several models: (1) UNet trained on only 1 datapoint and tested and validated on 2 additional datapoints, (2) a UNet trained on 73 datapoints (73-18-10 train-validation-test split), and (3) a DDUNet trained on the same 73 datapoints dataset, operating on $2\times 2$ subdomains. Furthermore, to assess the model’s sensitivity to random initialization, the training of the same architecture was repeated five times for the most relevant models. Based on these repetitions, the mean and standard deviation of the performance metrics were computed using the following equations:
@@
where $x_{i}$ denotes the metric value from the $i$-th training run, and $n=5$ is the number of runs. These results are summarized in Table 10 and Table 11. The choice of $n=5$ was made empirically to balance computational effort and statistical reliability. The standard deviations in  Tables 11 and 10 were used as validation: they are neither excessively large (indicating instability) nor unrealistically small (indicating insufficient sampling).
@@
In addition to testing the UNet and DDUNet trained on 73 datapoints on the 10-point test dataset, we also evaluate these models on the same datapoint used to test the UNet3DP.
Table 9: Performance metrics for predicting $T$ with different models and datasets. Errors in [C], MSE in [C], PATs in [%] and SSIM unitless. The LGCNN-test dataset corresponds to the 1 datapoint used for testing the LGCNN approach.
Model Data Case Huber L MAE MSE PAT SSIM
UNet3DP randomK3 train 0.0020 2.4954 0.0404 0.0040 6.43 0.8281
val 0.0269 5.2901 0.1365 0.0574 38.82 0.5717
test 0.0235 4.8642 0.1314 0.0492 39.05 0.5794
UNet rK101 train 0.0010 4.2443 0.0172 0.0021 2.34 0.9960
val 0.0051 4.1972 0.0441 0.0106 12.68 0.9859
test 0.0050 4.2140 0.0426 0.0104 11.90 0.9869
* LGCNN-test 0.0048 4.3985 0.0473 0.0100 13.63 0.9827
* scaling 0.0016 4.3426 0.0202 0.0033 4.33 0.9955
2$\times$2-DDUNet rK101 train 0.0018 3.8100 0.0236 0.0038 4.47 0.9940
val 0.0079 4.0084 0.0550 0.0165 17.03 0.9825
test 0.0076 3.7006 0.0549 0.0159 17.14 0.9835
* LGCNN-test 0.0063 3.4257 0.0548 0.0128 17.42 0.9804
* scaling 0.0025 4.1806 0.0235 0.0052 6.11 0.9940
@@
Table 10: Statistics for predicting $T$ with different models and datasets. Errors in [C], MSE in [C], PATs in [%] and SSIM unitless. The LGCNN-test dataset corresponds to the 1 datapoint used for testing the LGCNN approach. Mean $\pm$ standard deviation reported.
Model Data Case Huber L MAE MSE PAT SSIM
UNet3dp randomK3 train $0.0025\pm 0.0024$ $4.69\pm 0.17$ $0.0155\pm 0.0071$ $0.0063\pm 0.0062$ $1.92\pm 1.16$ $0.981\pm 0.012$
val $0.0219\pm 0.0008$ $4.89\pm 0.07$ $0.1114\pm 0.0023$ $0.0474\pm 0.0019$ $34.76\pm 0.99$ $0.696\pm 0.006$
test $0.0176\pm 0.0006$ $4.85\pm 0.09$ $0.1043\pm 0.0018$ $0.0368\pm 0.0015$ $34.76\pm 0.76$ $0.703\pm 0.006$
UNet101dp rK101 train $0.0011\pm 0.0002$ $4.37\pm 0.20$ $0.0182\pm 0.0020$ $0.0023\pm 0.0004$ $2.45\pm 0.27$ $0.995\pm 0.002$
val $0.0055\pm 0.0003$ $4.36\pm 0.16$ $0.0454\pm 0.0019$ $0.0114\pm 0.0006$ $12.96\pm 0.55$ $0.984\pm 0.002$
test $0.0052\pm 0.0003$ $4.35\pm 0.23$ $0.0441\pm 0.0019$ $0.0110\pm 0.0006$ $12.49\pm 0.48$ $0.985\pm 0.002$
* LGCNN-test $0.0049\pm 0.0002$ $4.30\pm 0.16$ $0.0470\pm 0.0010$ $0.0102\pm 0.0004$ $13.51\pm 0.59$ $0.983\pm 0.002$
* scaling $0.0017\pm 0.0001$ $4.48\pm 0.15$ $0.0208\pm 0.0014$ $0.0035\pm 0.0002$ $4.38\pm 0.17$ $0.995\pm 0.001$
2$\times$2-DDUNet101dp rK101 train $0.0014\pm 0.0003$ $4.11\pm 0.25$ $0.0203\pm 0.0026$ $0.00300\pm 0.0007$ $3.24\pm 0.68$ $0.995\pm 0.001$
val $0.0079\pm 0.0002$ $4.20\pm 0.25$ $0.0564\pm 0.0008$ $0.01648\pm 0.0005$ $17.32\pm 0.22$ $0.981\pm 0.002$
test $0.0075\pm 0.0001$ $4.05\pm 0.22$ $0.0552\pm 0.0008$ $0.01580\pm 0.0002$ $16.94\pm 0.36$ $0.982\pm 0.001$
* LGCNN-test $0.0057\pm 0.0003$ $4.00\pm 0.20$ $0.0526\pm 0.0015$ $0.01171\pm 0.0006$ $16.44\pm 0.63$ $0.981\pm 0.002$
* scaling $0.0025\pm 0.0001$ $4.04\pm 0.20$ $0.0251\pm 0.0007$ $0.00514\pm 0.0002$ $6.39\pm 0.17$ $0.994\pm 0.001$
LGCNN randomK3 Step 3 train $0.0001\pm 0.0001$ $2.64\pm 0.18$ $0.0064\pm 0.0008$ $0.0003\pm 0.0001$ $0.29\pm 0.08$ $0.996\pm 0.001$
val $0.0032\pm 0.0001$ $2.49\pm 0.29$ $0.0413\pm 0.0008$ $0.0065\pm 0.0003$ $10.88\pm 0.35$ $0.912\pm 0.003$
test $0.0025\pm 0.0000$ $2.74\pm 0.33$ $0.0382\pm 0.0006$ $0.0049\pm 0.0001$ $9.44\pm 0.35$ $0.918\pm 0.003$
scaling $0.0008\pm 0.0000$ $3.03\pm 0.31$ $0.0179\pm 0.0006$ $0.0016\pm 0.0001$ $2.54\pm 0.14$ $0.946\pm 0.005$

### v2

@@ Acknowledgements
Supplementary Material
[…]
Raw datasets:
  • Dataset of random permeability with 3+1 datapoints (3dp + 1)
  • Dataset of random permeability with 101 datapoints (101dp)
  • Dataset of real permeability with 4+1 datapoints (4dp + 1)
  • Prepared dataset for training the FNO, because the other datapoints are too large
Trained models (including hyperparameters):
[…]
  • LGCNN on random permeability fields, 3dp
[…]
  • LGCNN on real permeability fields, 4dp
  • vanilla models first trying to overfit them on 1 data point: PI-CNN, FNO
Code (including training and evaluation routines):
[…]
All supplementary material can be accessed via SURFDrive
under this link: https://surfdrive.surf.nl/files/index.php/s/f3Oqg3ufir9T9LL.
@@ A.1 Simulation Setup and Transport Regime
In Section˜2, we describe how the two datasets are generated using the subsurface simulation software Pflotran (Lichtner et al., 2015a), which solves the coupled mass and energy conservation equations. Here, we provide additional technical details and modeling assumptions relevant for reproducibility. Furthermore, we show that for the selected parameters, the heat transport in our system is dominated by advection by a theoretical analysis.
@@
Our data was generated on $\times12.8km$$\times$5m with a cell size of $\times5m$$\times$5m for the baseline simulations. To test scalability, we also simulate a domain that is twice as large in both $x$- and $y$-dimension for the baseline dataset, but only twice as large in $y$-dimension for the more realistic dataset due to dimension restrictions of the available input data.
The baseline dataset uses a constant hydraulic pressure gradient $\nabla p$ of 0.003 (Pelzer & Schulte, 2024; geo, 2022); permeability field $k$ is generated using Perlin noise (Perlin, 1985) within (1.02E-11, 5.10E-9) m2, and $100$ randomly placed heat pumps, which are all operating with a 5 °C injection temperature difference compared to the surrounding groundwater and an injection rate of 0.00024 m3/s. All values are based on realistic parameter ranges in the region of Munich(geo, 2022; Zosseder et al., 2022).
For the realistic dataset, instead of random permeability fields $k$, we use permeability field data that is derived from borehole measurements in the Munich region(Bayerisches Landesamt für Umwelt, 2015). Furthermore, based on subsurface measurements, we set $\nabla p=0.0025$ for these simulations. All other simulation parameters remain identical to the baseline. Other subsurface and operational parameters are taken directly from Pelzer & Schulte (2024), which also includes additional information about solver setup and boundaries. For mathematical details of the governing equations, we refer the reader to Lichtner et al. (2015b); Anderson (2007); Delleur (2016).
Notation Simplification
@@
To quantify whether heat transport in our system is dominated by advection or diffusion, we compute the dimensionless Péclet number (Rapp, 2017), which is defined as
@@
$\alpha=\frac{\kappa}{\rho c_{p}},$
with $\kappa$ the thermal conductivity, $\rho$ the density, and $c_{p}$ the specific heat capacity. We take the parameters of the solid phase of our simulation inputs of
  • Thermal conductivity: $\kappa=0.65\,\mathrm{W/(m\cdot K)}$
@@
With a characteristic length determined by the heat plume length of $2000\text{--}6000\,\mathrm{m}$ and mean x-velocity, we get:
@@
Reasoning for Decoupling Assumptions
In LGCNN, the goal is to split properties into local and global. To build separate datasets for the isolated steps, we require reasonable inputs and labels. Hence, the from data or simulation generated physical properties are analyzed as potential interim inputs / labels. From simulation, we get $\vec{v}$ (mostly local) and $T$ (heat is transported in an advective and diffusive manner $\rightarrow$ local and global). To split the temperature field into local and global, we handcraft a new property streamlines $\vec{s}$ that describe the central evolution of a heat plume driven only by advection, starting at each heat pump location. An important assumption underlying our approach is that the coupling between velocity and temperature can predominantly be treated as one-way in our setting. The velocity field strongly influences the temperature distribution but the temperature also influences the velocity. Since this effect mostly occurs in 3D domains with very high temperature gradients and we model 2D and very small temperature gradients, our assumption holds.
The general background velocity field influences the streamlines directly but the streamlines or rather the injection points and their mass influx also disturbs the local velocity field which in turn can influence other streamlines passing by from further upstream. By including the start locations of the streamlines, i.e., the locations of the heat pumps, as inputs to $\vec{v}$ and taking the steady-state flow field with running heat pumps from the simulation as a label, the influence of $\vec{s}$ is reduced and again decoupled as an additional input $i$ with the influence only local and already included in the labels. Hence, a one-way coupling from positions of heat pumps $i\rightarrow\vec{v}\rightarrow\vec{s}$ is possible.
A.2 NNs, Hyperparameters and Training Details
This section provides implementation and training details for all NN models used in this work. We begin with a description of the baseline UNet architecture, which also forms the foundation of both LGCNN and DDU-Net. Then, we outline the hyperparameter optimization process and present the final configurations per model and dataset. All models are trained using PyTorch (Paszke et al., 2017).
A.2.1 UNet Architecture
The UNet architecture used in this work is based on the original design presented by Ronneberger et al. (2015b), with several modifications to tailor it to the specific needs of our task of predicting temperature fields. In this section, we introduce the key concepts that define the architecture and explain how they contribute to the model’s design. These concepts will be employed in the hyperparameter search to identify the most suitable configuration, taking into account memory and data limitations.
@@ A.2.1 UNet Architecture
The training process also involves several hyperparameters. The learning rate controls how quickly the model adjusts its parameters during training. The weight decay parameter helps prevent overfitting by penalizing large weights in the model. Furthermore, the Adam optimizer (Kingma, 2014) is employed as optimizer. For the realistic permeability field, we additionally introduce the optimizable hyperparameter optimizer switch, which, when enabled, switches the optimizer to limited-memory BFGS (LBFGS) after 90 epochs.
During inference, each datapoint is processed as a whole, but during training, they are loaded in smaller patches cut out from the datapoint. Optimized hyperparameters include their overlap, i.e., inverse of skip per direction, and the patch size, called box length. The data split is untouched by this, i.e., separate simulation runs for train / val / test. This is important to not overlap test patches with training ones.
A.2.2 Energy-Based Loss for PINNs
For our PINN, we overfitted a UNet with an energy loss $\mathcal{L}_{\text{energy}}$ to a single datapoint from 3dp.
The physics-informed loss for predicting the temperature field $T$ considers energy conservation, including advective and conductive terms, and heat pump contributions. Let $\vec{v}=(v_{x},v_{y})$ be the velocity field and $p$ the pressure. Denote the predicted temperature as $\hat{T}$. The energy loss is defined as
$\mathcal{L}_{\text{energy}}=\text{MSE}\Big(E-E^{\text{HP}}_{\text{conv}},\,0\Big),$ (5)
where $E$ is the local energy residual and $E^{\text{HP}}$ the energy inflow from heat pumps.
Energy residual.
The energy residual accounts for advection and thermal diffusion:
$\rho_{m}$ $=\text{molar density}(\hat{T},p),$ (6)
$h$ $=\text{enthalpy}(\hat{T},p),$ (7)
$E_{u}$ $=\frac{\partial}{\partial x}\Big(\rho_{m}\,h\,v_{x}-\frac{\kappa}{\Delta x}\frac{\partial\hat{T}}{\partial x}\Big),$ (8)
$E_{v}$ $=\frac{\partial}{\partial y}\Big(\rho_{m}\,h\,v_{y}-\frac{\kappa}{\Delta x}\frac{\partial\hat{T}}{\partial y}\Big),$ (9)
$E$ $=E_{u}+E_{v},$ (10)
where $\kappa$ is the thermal conductivity and $\Delta x$ is the uniform grid spacing.
Heat pump contribution.
Heat pumps add localized energy sources based on a one-hot-encoding of the positions of heat pumps (ID) to get a spatially distributed contribution:
$\rho$ $=\text{density}(\hat{T},p),$ (11)
$E^{\text{HP}}$ $=c_{p}\,\rho\,\Delta T\,Q\,\frac{\text{ID}}{\Delta x^{3}},$ (12)
$E^{\text{HP}}_{\text{conv}}$ $=K*E^{\text{HP}},$ (13)
where $c_{p}$ is the specific heat capacity of water, $\Delta T$ is the temperature difference at the injection well, $Q$ is the volumetric flow rate, ID is the material indicator (1 if heat pump present, 0 otherwise), and $K$ is a fixed, handcrafted $3\times 3$-convolutional kernel $K=\begin{bmatrix}-1&0&1\\ 0&0&0\\ 1&0&-1\end{bmatrix}$ with zero-padding used to smooth the inflow contribution. It distributes the inflow over neighboring cells while preserving directional structure and its shape matches our observations.
A.2.3 Hyperparameters and Hyperparameter Optimization
We optimize the hyperparameters of our architecture, training process and data loading in several rounds with an automated tree-based search using Optuna (Akiba et al., 2019) and additional manual tweaking. Optuna performs optimization using a tree-structured Parzen estimator algorithm. An overlook of all considered hyperparameters, their ranges and our final choice, as well as the hyperparameters fixed during optimization is provided per used model architecture and dataset, e.g., purely data-driven approaches, LGCNN on individual steps or the full pipeline, on datasets of random $k$ versus realistic $k$.
@@ A.2.3 Hyperparameters and Hyperparameter Optimization
The inputs to the model are pkipki (pressure field, permeability field, and location of heat pumps). For training, the MAE loss function was used, and optimized with the Adam optimizer. The learning rate is set to $1\times 10^{-5}$. The model is trained for 10,000 epochs.
Hyperparameters for UNet variants on 101dp
The values considered during hyperparameter optimization for the UNet and DDUNet, trained on 73 datapoints and performed using Optuna (Akiba et al., 2019), along with the best settings found, are listed in Table˜2. Certain hyperparameters were fixed: we set the number of epochs to 750, with an early stopping criterion based on validation loss and a patience of 80 epochs. Additionally, we note that some hyperparameter combinations (e.g., 32 initial features, depth 6, and 3 convolutions per layer with a kernel size of 7) caused memory issues, leading to their exclusion from the hyperparameter search.
@@
Hyperparameter Range $pki\rightarrow v_{x}v_{y}$ (Step 1) $iv_{x}v_{y}ss_{o}k\rightarrow T$ (Step 3) $pki\rightarrow T$ (Full)
@@
After the hyperparameter search, the values corresponding to the best-performing model (based on Huber loss for the validation dataset) were selected. With these values fixed, five models were trained using different randomly sampled initializations to evaluate sensitivity to random initialization, for these values see Tables˜8 and 7.
Hyperparameters for UNet+++ and dilated UNet
[…]
Table 3: Baseline hyperparameters for end-to-end temperature prediction.
Parameter Dilated CNN (3dp) UNet++ (101dp)
Parameters of the Dataset
dataset size 3 101
inputs pikpik ikik
cutouts / patches full domain full domain
Parameters of Training
batch size 20 2
training loss MAE MSE
initial learning rate $10^{-4}$ (manual decay) $2\times 10^{-4}$
optimizer Adam Adam
epochs 1000 190
Parameters of the Network
architecture Dilated CNN (zero-padded) UNet++
kernel size 5 default
depth 3 default
initial features 32 default
stride 1 default
dilation 10 default
activation function ReLU default
normalization BatchNorm default
repeat inner blocks False -
Hyperparameters for PINN and FNO
Hyperparameters of our PI-CNN and FNO trainings are listed in Table˜4. For PI-CNN, we used the optimized, well-working architecture from before; for FNO a transferred set of hyperparameters with an additional hyperparameter search on layers, modes, inputs and optimizer.
Table 4: Baseline hyperparameters for overfitting experiments on a single datapoint (dp id 0).
Parameter FNO (overfit test) PINN (overfit test)
Parameters of the Dataset
dataset size 1 1
inputs pikpik $pikv_{x}v_{y}$
box length 480 256
Parameters of Training
batch size 8 20
training loss MSE MSE (physics-only)
initial learning rate $10^{-3}$ (LBFGS) $10^{-4}$ (manual decay)
optimizer LBFGS Adam
epochs 1000 500
Parameters of the Network
architecture FNO UNet (zero-padding)
modes / kernel size 64 modes $4\times 4$
depth 6 4
initial features 16 32
stride 1 1
dilation 1 1
activation function ReLU ReLU
normalization BatchNorm BatchNorm
repeat inner blocks False False
Hyperparameters for LGCNN on random permeability
The values considered during hyperparameter optimization with Optuna (Akiba et al., 2019) and the best settings found for both steps of LGCNN are listed in Table˜5. Although the optimization was originally run for 100 epochs, the optimum was consistently found within the first 25 epochs. Therefore, to reduce computation cost, we therefore conservatively lowered the maximum number of epochs to 50. This adjustment does not affect any of the reported metrics in the paper.
@@
Parameter Range Step 1 ($v$) Step 3 ($T$)
@@
depth 1, 2, 3, 4 4 4
Fixed parameters for this hyperparameter search are the learning rate (fixed at $10^{-4}$), ReLU as activation function, the batch size of 20, and the use of a batch normalization layer within the inner blocks of the UNet architecture. The validation loss used for selecting the optimal model is the MAE.
Hyperparameters for LGCNN on real permeability fields
The values considered during hyperparameter optimization on the dataset with a more realistic permeability field were selected using Optuna (Akiba et al., 2019), and are summarized in  Table˜6, along with the best configurations found for both steps of the LGCNN. The optimization was run for up to 100 epochs. For more background on the network architecture and the various hyperparameters, cf. A.2.1. Several hyperparameters were fixed during this process. These include a constant learning rate schedule, an Adam optimizer with a weight decay of $10^{-4}$, and, when enabled, a switch to LBFGS after 90 epochs. Fixed architectural parameters include a convolutional stride and dilation of 1. During training, the inputs were cut out from the full datapoints. For model comparison, the validation loss was consistently computed using the Huber loss.
@@
No. initial features 8, 16, 32 16 16
kernel size 3, 5 5 5
depth 4, 5, 6 6 6
@@ A.3 Additional Experimental Results
This section provides additional experimental results. While the main results section focused only on the test and scaling datasets, we also include here the metric values on training and validation datasets. For easier comparison, the test and scaling metrics are re-listed as well.
Vanilla UNet and DDUNet on 101dp: Metrics of training and ablation study
We present metrics for predicting the temperature field directly from the inputs pkipki using a data-driven approach, evaluated on the training, validation, and test datasets. The results are provided for two models: (1) UNet trained on 73 datapoints (73-18-10 train-validation-test split), and (2) DDUNet trained on the same 73 datapoints dataset, operating on $2\times 2$ subdomains. To assess the model’s sensitivity to random initialization, the training of the same architecture was repeated five times. Based on these repetitions, the mean and standard deviation of the performance metrics were computed using the following equations:
@@
where $x_{i}$ denotes the metric value from the $i$-th training run, and $n=5$ is the number of runs. These results are summarized in Table˜7. The choice of $n=5$ was made empirically to balance computational effort and statistical reliability. The standard deviations in  Table˜7 were used as validation: they are neither excessively large (indicating instability) nor unrealistically small (indicating insufficient sampling).
@@
In addition to testing the UNet and DDUNet trained on 73 datapoints on the 101dp test dataset, we also evaluate these models on the same datapoint used to test the UNet3dp.
@@
Table 7: Statistical error metrics for end-to-end prediction of $T$ with 101dp UNet and DDUNet. Errors in [°C], MSE in [°C2], and SSIM unitless. The LGCNN-test dataset corresponds to the 1 datapoint used for testing the LGCNN approach. Mean $\pm$ standard deviation reported.
Model Data Case Huber L MAE MSE
UNet 101dp train $0.0011\pm 0.0002$ $4.37\pm 0.20$ $0182\pm 0.0020 0.0$$0023\pm 0.0004$
val $0.0055\pm 0.0003$ $4.36\pm 0.16$ $0454\pm 0.0019 0.0$$0114\pm 0.0006$
test $0.0052\pm 0.0003$ $4.35\pm 0.23$ $0441\pm 0.0019 0.0$$0110\pm 0.0006$
* LGCNN-test $0.0049\pm 0.0002$ $4.30\pm 0.16$ $0470\pm 0.0010 0.0$$0102\pm 0.0004$
* scaling $0.0017\pm 0.0001$ $4.48\pm 0.15$ $0208\pm 0.0014 0.0$$0035\pm 0.0002$
DDUNet 101dp train $0.0014\pm 0.0003$ $4.11\pm 0.25$ $0203\pm 0.0026 0.0$$00300\pm 0.0007$
val $0.0079\pm 0.0002$ $4.20\pm 0.25$ $0564\pm 0.0008 0.0$$01648\pm 0.0005$
test $0.0075\pm 0.0001$ $4.05\pm 0.22$ $0552\pm 0.0008 0.0$$01580\pm 0.0002$
* LGCNN-test $0.0057\pm 0.0003$ $4.00\pm 0.20$ $0526\pm 0.0015 0.0$$01171\pm 0.0006$
* scaling $0.0025\pm 0.0001$ $4.04\pm 0.20$ $0251\pm 0.0007 0.0$$00514\pm 0.0002$
Vanilla UNet on 3dp: Metrics of training and ablation study
We present metrics for predicting $T$ directly from the inputs pkipki with UNet3dp in Table˜8, evaluated on training, validation, test and scaling datapoints. To assess the model’s sensitivity to random initialization, the training of the same architecture was repeated five times. Other models were not trained again to reduce computational effort and because the results of all vanilla approaches indicate that the performance of UNet101dp, UNet3dp, and DDUNet101dp can be viewed as a representative of the general performance of purely data-driven methods.